# Deconvoluting Thermomechanical Effects in X-ray Diffraction Data using Machine Learning

### v1

@@ Abstract
X-ray diffraction is ideal for probing sub-surface state during complex or rapid thermomechanical loading of crystalline materials. However, challenges arise as the size of diffraction volumes increase due to spatial broadening and inability to deconvolute the effects of different lattice deformation mechanisms. Here we present a novel approach to use combinations of physics-based modeling and machine learning to deconvolve thermal and mechanical elastic strains for diffraction data analysis. The method builds on a previous effort to extract thermal strain distribution information from diffraction data. The new approach is applied to extract the evolution of thermomechanical state during laser melting of an Inconel 625 wall specimen which produces significant residual stress upon cooling. A combination of heat transfer and fluid flow, elasto-plasticity, and X-ray diffraction simulations are used to generate training data for machine-learning (Gaussian Process Regression, GPR) models that map diffracted intensity distributions to underlying thermomechanical strain fields. First-principles density functional theory is used to determine accurate temperature-dependent thermal expansion and elastic stiffness used in the elasto-plasticity modeling. The trained GPR models are found to be capable of deconvoluting the effects of thermal and mechanical strains, in addition to providing information about underlying strain distributions, even from complex diffraction patterns with irregularly shaped peaks.
@@ 1 Introduction
Understanding and controlling the development of residual stress during traditional welding and now, additive manufacturing (AM), is an on-going challenge. Rapid, localized heating then solidification leads to large thermal stresses which, in turn, lead to plastic flow and residual stress. These residual stresses can drive hot-cracking or cause such severe distortions that a part is unusable. Optical and thermography measurements provide a means to characterize the temperature gradients driving stress development [1, 2, 3, 4, 5, 6, 7, 8], but are limited to the surface. In addition, while there has been a significant focus placed on developing high-speed X-ray imaging at synchrotron sources to monitor melt pool characteristics and the development of porosity, in support of AM processing, the use of scattering and the study of stress development has been, in comparison, limited. However, this is changing and recent efforts are building on upon imaging efforts to also use scattering to study the complex rapid solidification processes [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. More specifically, diffraction processes are capable of monitoring the development of texture, phases, and residual strains (and related stress).
@@
An acute challenge in using diffraction to monitor stress development in combined thermal and mechanical loading is that both temperature and mechanical stress alter the state of the crystal lattice (which is what is probed during diffraction measurements). Specifically in cubic crystal types, both temperature and hydrostatic stress will alter the volumetric portion of the lattice strain tensor in the same fashion, making deconvolution of the effects difficult. Assumptions can be made regarding whether the thermal or mechanical alterations to lattice state are dominant for analysis, but these can introduce significant uncertainty into interpretation of data when magnitudes become approximately equal. In addition, the spread of thermal and mechanical strain distributions through a diffraction volume can be extensive, particularly when using high-energy X-rays in transmission through thick specimens. A single diffraction image (projection), does not provide sufficient information for direct reconstruction of the distributions present.
@@
To overcome these challenges, we present a machine learning (ML) approach in which Gaussian Process Regression (GPR) models are trained to learn the relationship between diffraction patterns and underlying temperature and mechanical strain distributions. The ML training process is supported by physics-based modeling of the heating and cooling processes that leads to stress development, along with accurate diffraction simulations. This modeling, along with diffraction simulation, is used to create a ‘dictionary’ of diffraction patterns for ML model training. The trained ML model is then transferred to the analysis of experimental data. To attain accurate thermal and mechanical properties, density functional theory (DFT) is utilized. This work builds on our previous effort to extract temperature (thermal strain) distribution metrics from in situ diffraction patterns [20]. However, as mentioned, due to convolution of the effects of thermal and mechanical strain, this previous effort was only accurate at high-temperatures. Transfer learning is particularly appropriate for cases where the underlying physics is fairly well understood and modeled. Any uncertainty in the accuracy of the model may have implications to the ”transferring” to real data. Thus, high fidelity models are required and the improvements to our previous work have been made through the following advances i) the use of DFT for temperature-dependent thermomechanical property, ii) modeling of stress distributions using a thermomechanical elasto-plasticity model, iii) update of the X-ray diffraction modeling framework to incorporate anisotropic lattice strains during mechanical loading, and iv) update of the GPR model training to utilize anisotropic diffraction ring expansion and contraction.
@@
The approach is demonstrated through extraction of thermal and elastic strain distribution metrics from a wall specimen of Inconel 625 specimen in a transmission geometry that has become a standard for in situ AM experiments. An overview of the data-processing workflow is provided in Figure 1. The manuscript begins by describing the experimental data to be analyzed in §2. Next in §3, summaries of the various physics models used for GPR model training are provided in §4 . The GPR models used and their training is next described. The trained GPR models are then applied to experimental in situ diffraction data in §5 . Our approach and avenues for future study are discussed in §6
@@
Figure 1: Overview of the simulations used to train Gaussian Process Regression (GPR) models for X-ray diffraction data analysis. The trained models are subsequently transferred to the target (experimental) domain to separate thermal and mechanical strain effects from experimental X-ray diffraction data. Temperature fields $T(\bm{r},t)$ and temperature-dependent properties (Young’s modulus $E(T)$, and Poisson’s ratio $\nu(T)$, coefficient of thermal expansion $\alpha(T)$ are used as input for thermomechanical modeling to predict evolving thermal $\varepsilon^{T}(\bm{r},t)$ and elastic strains $\bm{\varepsilon}^{E}(\bm{r},t)$. These strains are used for input into X-ray diffraction simulations and GPR model training. The trained GPR models are then used for analysis of experimental data.
@@ 2 Material and Experiment Description
The material tested was Inconel 625 (IN625) that was extracted from a block built using laser powder bed fusion at the National Institute of Standards and Technology (NIST) [21]. The block was built in an EOS M290 system using powder also obtained from the EOS. The block dimensions were 50 mm $\times$ 15 mm $\times$ 6 mm where the 6 mm direction is the build direction. The build followed manufacturer recommendations [22] of laser power of 285 W, laser speed of 960 mm/s, and interlayer rotation of 67.5. After build, the specimen was stress-relief heat treated at 800 C for 2 hr. A wall specimen was then extracted using electro-discharge machining with dimensions of 15 mm $\times$ 0.53 mm $\times$ 3 mm, with the 3 mm dimension being aligned with the build dimension. Previous characterization of the material found that the grain size ranged from approximately 10 $\mu$m to 100 $\mu$m [20].
The wall specimen was then used for an in situ laser melting experiment at Sector 1-ID of the Advanced Photon Source. An existing in operando laser powder bed fusion (LPBF) simulator [23] consisting of a laser and control, sealed chamber, and sample staging was used for laser melting. A schematic of the experimental geometry is provided in Fig. 2. The laser system was composed of a ytterbium fiber laser (IPG YLR-500-AC) controlled with an intelliSCANde 30. Using this system, the laser was rastered along the top of the wall specimen in the $\bm{x}$ direction at velocity $\bm{v}$ with magnitude of 0.05 m/s and power $P$ of 120 W (2,400 J/m), generating a relatively large high-temperature region and resulting residual stresses. During laser melting, a primarily uniaxial tensile residual stress developed along the top of the wall specimen. This process consisted of i) rapid expansion during heating, ii) tensile compressive stresses sufficiently high to generate plastic flow and permanent compression, and iii) mechanical extension and tensile loading as the specimen cooled to accommodate compatibility. The primary tensile mechanical ($\varepsilon^{E}_{xx}$ strains that develop during cooling are also illustrated on the specimen in Fig 2.
Figure 2: Schematic of the (simulation and experiment) geometry for the wall laser melting experiment.
@@
Figure 3: a) Illustration of the four sets of azimuthal bins (colored) around which diffraction images were integrated. b) Concatenation of the 1D intensity data from the different bins for use in ML models. A comparison of representative experimental and simulated diffraction data are shown.
@@ 3 Training Data Generation
3.1 First-Principles Thermomechanical Property Calculations
The literature is limited regarding temperature-dependent properties of specific alloy compositions. Here we utilized first-principles DFT using the VASP code [24] to generate temperature-dependent IN625 elastic moduli and coefficients of thermal expansion for input into later thermomechanical modeling. Our process is described in detail in [25, 26], but is summarized here. In this approach, the Helmholtz energy $F$ has a 0 K static energy contribution, a vibrational contribution, and a thermal electron contribution. Here, each individual contribution to the Helmholtz energy is predicted independently, with the vibrational and thermal electron contributions calculated as a function of volume $V$ at a given temperature $T$, and then the three contributions summed. From the temperature-dependent Helmholtz energy , the equilibrium volume $V^{0}$ is calculated as a function of temperature. The linear coefficient of thermal expansion $\alpha$ is then calculated as
$\alpha(T)=\frac{1}{3V^{0}}\left(\frac{\partial V^{0}}{\partial T}\right)|_{P}\quad.$ (1)
For the IN625 property modeling, we used the compositions reported for AM Bench 2018 material [21] of which the same powder used was used here. In the IN625, the major alloying components besides Ni are Cr (20.61 wt%), Nb (3.97 wt%), and Mo (8.82 wt%). Based on CALPHAD predictions, there are two major phases, i.e., the dominant FCC matrix ($\gamma$) and the topologically close-packed (TCP) phase, $\delta$. Here we assume that the thermal expansion and elastic properties are primarily dependent on the FCC matrix. The composition of the FCC matrix according to CALPHAD predictions is roughly Ni21Cr8Mo2Nb1 [26]. To predict properties, relatively small cubic supercells (32-atom) were modeled and then the results averaged to strike a balance between accuracy and computational efficiency. We generated 6 special quasirandom structures (SQSs) [27] and 6 supercells in random approximates (SCRAPs) [28]. Figure 4 shows two representative SQS and SCRAPs structures, with Fig. 4a showing one of the generated SQSs and Fig. 4b one of the generated SCRAPs. Temperature-dependent Helmholtz energy was calculated for each of these supercells and then used for both coefficient of thermal expansion and elastic moduli calculations, see details including first-principles settings in [26]. Figure 5a shows the calculated temperature-dependent coefficient of thermal expansion $\alpha(T)$ using this procedure. The values matched well with experimental data used in [20] within the temperature bounds measured.
Figure 4: A One of the generated SQSs and B) one of the generated SCRAPs of Ni21Cr8Mo2Nb1 (32-atom supercell), where the colored spheres represent different atom types: Ni (gray), Cr (blue), Mo (purple), and Nb (green).
Single crystal elastic moduli ($C_{ij}$) as a function of temperature $T$ were predicted using the quasi-static approach (QSA) [29, 25, 30] according to the predicted relationship between $C_{ij}$ and $V$ at 0 K and the previously calculated relationship between $T$ and $V$ (i.e., thermal expansion) [26], resulting in the predicted $C_{ij}(T)$. Note that the DFT-based calculations of $C_{ij}(V)$ at 0 K were predicted using the anisotropic form of Hooke’s law [31], where the employed non-zero strains applied are $\pm$ 0.01. Since both the SQSs and SCRAPs of IN625 are monoclinic configurations, averaged values of the elastic moduli were used to calculate cubic elastic moduli $\bar{C}_{ij}$, i.e., $\bar{C}_{11}=(C_{11}+C_{22}+C_{33})/3$, $\bar{C}_{12}=(C_{12}+C_{13}+C_{23})/3$, and $\bar{C}_{44}=(C_{44}+C_{55}+C_{66})/3$. Figure 5b shows the cubic single crystal moduli $\bar{C}_{11}$, $\bar{C}_{12}$, and $\bar{C}_{44}$ determined from the DFT calculations. From the figure, we see that both $\bar{C}_{11}$ and $\bar{C}_{12}$ decrease faster than $\bar{C}_{44}$ with increasing temperature.
Figure 5: Temperature-dependent a) coefficient of thermal expansion $\alpha$ and b) cubic single crystal elastic moduli $\bar{C}_{ij}$ of AM IN625 calculated using DFT-based quasi-harmonic analysis (QHA) and quasi-static analysis (QSA) respectively.
For input into the continuum thermomechanical model (§3.3), $\bar{C}_{11}$, $\bar{C}_{12}$, and $\bar{C}_{44}$ were used to generate isotropic elastic moduli. The mean of the Reuss and Voigt averages of an untextured cubic polycrystal [32] were used to calculate Young’s Modulus $E(T)$ and the shear modulus $G(T)$, from which the Poisson’s ratio $\nu(T)$ was calculated. For temperatures beyond the melting temperature of the material, the bulk modulus of the material was assumed to remain constant while the shear modulus was reduced to 5% of the value prior to melting. The resulting temperature dependent elastic moduli used for thermomechanical model input are provided in Fig. 6.
Figure 6: Temperature-dependent isotropic elastic moduli (a) Young’s modulus $E$ and b) Poisson’s ratio $\nu$) used as input for the finite element simulations of the development of thermal and residual stresses in the AM IN625 wall specimens derived from the single crystal elastic moduli calculated from DFT shown in Fig. 5b.
@@ 3.2 Heat Transfer and Fluid Flow
Specific Heat ($\mathrm{J}\mathrm{/}\mathrm{k}\mathrm{g}\mathrm{/}\mathrm{K}$) $360.4+0.26~{}T-4\times 10^{-5}~{}T^{2}$
@@
A series of simulations were performed rastering a simulated laser heat source over the top of wall specimens with the same thickness (0.53 mm) and height (3 mm) as in the experiment. The length of the specimen was 11.5 mm to reduce computational cost. The heat source moved along the top surface (normal $\bm{z}$) in the $\bm{x}$ direction. Following [20], nine different simulations were performed around the nominal experimental laser setting (power and speed varied) with the same approximate laser size and are summarized in Table 2. After the heat source passed over the length of the specimen, the sample was allowed to cool for 5 s as mentioned. The primary output are temperature fields $T(\bm{r})$, where $\bm{r}$ denotes position, output at a rate of 500 Hz as the heat source moved and then 100 Hz as the sample cooled for a total of 150 time steps. As mentioned, the code adaptively changes the calculation grid throughout the simulation for computational efficiency. To prepare the temperature fields for use in the next steps of the workflow (Fig. 1), the simulations were post-processed using bi-linear interpolation to remap the output regular grid with a point spacing of 20 $\mu$m.
@@ 3.3 Thermomechanical Model
Once the 9 time series temperature fields were calculated from the heat-transfer and fluid flow modeling, they are used as external data inputs for an elasto-plasticity model in ANSYS to generate thermal $\varepsilon^{T}(\bm{r})$ and elastic strain $\bm{\varepsilon}^{E}(\bm{r})$ fields. The element type used was eight node ‘brick’ elements with three degrees of freedom (translations in $\bm{x}$, $\bm{y}$, and $\bm{z}$) at each node. Element sizes ranged from 125 to 140 $\mu$m (approximately 10,000 elements per simulations) along their edges.
@@
where $\sigma^{Y}(T)$ is the initial temperature-dependent yield strength, $H(T)$ is the temperature-dependent hardening rate, $\tilde{\varepsilon}^{P}$ is the equivalent plastic strain, and $\tilde{\sigma}$ is the equivalent (von Mises) stress. If the stresses are sufficient to initiate yielding (and ultimately the development of residual stress), plastic flow occurs. The elastic moduli (from DFT) used for Eq. 4 were provided in Fig. 6. The tabulated yield and hardening parameters were determined from a combination of mechanical tests ($\leq 500^{\circ}$) performed on the same material and IN625 parameters built into the ANSYS package (Table 3).
Each time step takes the previously calculated thermal field and calculates local thermal strains (Eq. 3) from the temperature-dependent coefficient of thermal expansion previously calculated using DFT [26]. A strain field is then calculated which satisfies both compatibility (in conjunction with the thermal and plastic strain) and mechanical equilibrium. Fig. 7 shows representative fields ($P=120$ W, $v=0.05$ m/s, and spot diameter of 100 $\mu$m) calculated by the thermomechanical model as the laser passes over the sample (Figs. 7a, 7c,7e, and 7g) and after cooling (Figs. 7b, 7d,7f, and 7h). The thermal and elastic strains are plotted with different color scales due to the large differences in maximum magnitude. Important to note is the relatively large residual tensile $\varepsilon^{E}_{xx}$ strains that develop during the cooling process.
@@
Figure 7: Representative fields from the thermomechanical simulations ($P=120$ W, $v=0.05$ m/s, and spot diameter of 100 $\mu$m) as the laser is passing overhead: a) $T$, c) $\varepsilon^{T}$, e) $\varepsilon^{E}_{xx}$, and g) $\varepsilon^{E}_{yy}$. Representative fields from the same laser parameter simulation after cooling: b) $T$, d) $\varepsilon^{T}$, f) $\varepsilon^{E}_{xx}$, and h) $\varepsilon^{E}_{yy}$.
To facilitate X-ray diffraction simulations from the elasto-plasticity results, a series of scripts were developed to convert ANSYS mechanical data into a format compatible with X-ray simulations. This effort included developing a new (i) Python based binding for ASCII output of ANSYS data; (ii) a script to convert element position, stress, and strain data into the coordinate system used by HEXRD; and (iii) developing a framework to interpolate thermomechanical data on to a finer grid (20 $\mu$m spacing) necessary for diffraction simulations.

### v3

@@ Abstract
X-ray diffraction is ideal for probing sub-surface state during complex or rapid thermomechanical loading of crystalline materials. However, challenges arise as the size of diffraction volumes increases due to spatial broadening and inability to deconvolute the effects of different lattice deformation mechanisms. Here, we present a novel approach to use combinations of physics-based modeling and machine learning to deconvolve thermal and mechanical elastic strains for diffraction data analysis. The method builds on a previous effort to extract thermal strain distribution information from diffraction data. The new approach is applied to extract the evolution of thermomechanical state during laser melting of an Inconel 625 wall specimen which produces significant residual stress upon cooling. A combination of heat transfer and fluid flow, elasto-plasticity, and X-ray diffraction simulations are used to generate training data for machine-learning (Gaussian Process Regression, GPR) models that map diffracted intensity distributions to underlying thermomechanical strain fields. First-principles density functional theory is used to determine accurate temperature-dependent thermal expansion and elastic stiffness used for elasto-plasticity modeling. The trained GPR models are found to be capable of deconvoluting the effects of thermal and mechanical strains, in addition to providing information about underlying strain distributions, even from complex diffraction patterns with irregularly shaped peaks.
@@ 1 Introduction
Understanding and controlling the development of residual stress during traditional welding, and now additive manufacturing (AM), is an ongoing challenge. Rapid, localized heating then solidification leads to large thermal stresses which, in turn, lead to plastic flow and residual stress. These residual stresses can drive hot-cracking or cause such severe distortions that a part is unusable. Optical and thermography measurements provide a means to characterize the temperature gradients driving stress development [1, 2, 3, 4, 5, 6, 7, 8], but are limited to the surface. In addition, while there has been a significant focus placed on developing high-speed X-ray imaging at synchrotron sources to monitor melt pool characteristics and the development of porosity in support of AM processing, the use of scattering and the study of stress development has been limited in comparison. However, recent efforts are building upon imaging efforts to also use scattering to study complex rapid solidification processes [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. More specifically, diffraction processes are capable of monitoring the development of texture, phases, and residual strains (and related stress).
@@
An acute challenge in using diffraction to monitor stress development in combined thermal and mechanical loading is that both temperature and mechanical stress alter the state of the crystal lattice (which is what is probed during diffraction measurements). Specifically in cubic crystal types, both temperature and hydrostatic stress will alter the volumetric portion of the lattice strain tensor in the same fashion, making deconvolution of the effects difficult. Assumptions can be made regarding whether the thermal or mechanical alterations to lattice state are dominant for analysis, but these can introduce significant uncertainty into interpretation of data when magnitudes become approximately equal. In addition, the spread of thermal and mechanical strain distributions through a diffraction volume can be extensive, particularly when using high-energy X-rays in transmission through thick specimens. A single diffraction image (projection) does not provide sufficient information for direct reconstruction of the distributions present.
@@
To overcome these challenges, we present a machine learning (ML) approach in which Gaussian Process Regression (GPR) models are trained to learn the relationship between diffraction patterns underlying thermomechanical strain distributions. The ML training process is supported by physics-based modeling of the heating and cooling processes that leads to stress development. This modeling, along with accurate diffraction simulations, is used to create a ‘dictionary’ of diffraction patterns for ML model training. The trained ML model is then transferred to the analysis of experimental data. To attain accurate thermal and mechanical properties, density functional theory (DFT) is utilized. This work builds on our previous effort to extract temperature (thermal strain) distribution metrics from in situ diffraction patterns [20]. However, as mentioned, due to convolution of the effects of thermal and mechanical strain, this previous effort was only accurate at high temperatures. Transfer learning is particularly appropriate for cases where the underlying physics is fairly well understood and modeled. Any uncertainty in the accuracy of the model may have implications to the “transferring” to real data. Thus, high fidelity models are required and the improvements to our previous work have been made through the following advances: i) the use of DFT for temperature-dependent thermomechanical properties, ii) modeling of stress distributions using a thermomechanical elasto-plasticity model, iii) update of the X-ray diffraction modeling framework to incorporate anisotropic lattice strains during mechanical loading, and iv) update of the GPR model training to utilize anisotropic diffraction ring expansion and contraction.
@@
The approach is demonstrated through extraction of thermal and elastic strain distribution metrics from a wall specimen of Inconel 625 in a transmission geometry that has become a standard for in situ AM experiments. An overview of the data-processing workflow is provided in Fig. 1. This work begins by describing the experimental data to be analyzed in §2. Next in §3, summaries of the various physics models used for GPR model training are provided. The GPR models used and their training is described in §4. The trained GPR models are then applied to experimental in situ diffraction data in §5. The work finishes by discussing our approach and avenues for future study in §6.
@@
Figure 1: Overview of the simulations used to train Gaussian Process Regression (GPR) models for X-ray diffraction data analysis. The trained models are subsequently transferred to the target (experimental) domain to separate thermal and mechanical strain effects from experimental X-ray diffraction data. Temperature fields $T(\bm{r},t)$ and temperature-dependent properties (Young’s modulus $E(T)$, Poisson’s ratio $\nu(T)$, and coefficient of thermal expansion $\alpha(T)$) are used as input for thermomechanical modeling to predict evolving thermal $\varepsilon^{T}(\bm{r},t)$ and elastic strains $\bm{\varepsilon}^{E}(\bm{r},t)$. These strains are used for input into X-ray diffraction simulations and GPR model training. The trained GPR models are then used for analysis of experimental data.
@@ 2 Material and Experiment Description
The material tested was Inconel 625 (IN625) that was extracted from a block built using laser powder bed fusion at the National Institute of Standards and Technology (NIST) [21]. The block was built in an EOS M290 system using powder also obtained from EOS. The block dimensions were 50 mm $\times$ 15 mm $\times$ 6 mm where the 6 mm direction is the build direction. The build followed manufacturer recommendations [22] of laser power of 285 W, laser speed of 960 mm/s, and interlayer rotation of 67.5. After build, the specimen was stress-relief heat treated at 800 C for 2 hr. A wall specimen was then extracted using electro-discharge machining with dimensions of 15 mm $\times$ 0.53 mm $\times$ 3 mm, with the 3 mm dimension being aligned with the build dimension. Previous characterization (including electron backscatter diffraction) of the material found that the grain size ranged from approximately 10 $\mu$m to 100 $\mu$m with a mean between 25 $\mu$m and 30 $\mu$m [20].
The wall specimen was then used for an in situ laser melting experiment at Sector 1-ID of the Advanced Photon Source. An existing in operando laser powder bed fusion (LPBF) simulator [23] consisting of a laser and control, sealed chamber, and sample staging was used for laser melting. A schematic of the experimental geometry is provided in Fig. 2. The laser system was composed of a ytterbium fiber laser (IPG YLR-500-AC) controlled with an intelliSCANde 30. Using this system, the laser was rastered along the top of the wall specimen in the $\bm{x}$ direction at velocity $\bm{v}$ with magnitude of 0.05 m/s and power $P$ of 120 W (2,400 J/m), generating a relatively large high-temperature region and resulting residual stresses. During laser melting, a primarily uniaxial tensile residual stress developed along the top of the wall specimen. This process consisted of i) rapid expansion during heating, ii) compressive stresses sufficiently high to generate plastic flow and permanent compression, and iii) extension and tensile loading as the specimen cooled to accommodate compatibility. The primary tensile mechanical ($\varepsilon^{E}_{xx}$) strains that develop during cooling are also illustrated on the specimen in Fig 2.
Figure 2: Schematic of the (simulation and experiment) geometry for the wall laser melting experiment. The inset indicates points where diffraction measurements and simulations were performed with respect to the top of the sample. The X-ray beam was centered 20 $\mu$m below the sample surface during the experiment (shaded green), while the X-ray beam was centered 20 $\mu$m, 40 $\mu$m, and 60 $\mu$m below the sample surface for various simulations used in Gaussian Process Regression model training.
@@
Figure 3: Representative diffraction images from a) the experimental sample before laser melting, b) the experimental sample after laser melting, c) a thermomechanical simulation used for Gaussian Process Regression surrogate model training. Maximum intensity thresholds have been selected for each image to make the diffraction peaks and their extent visible.
For use in ML models (described below), each diffraction image was then integrated into six different azimuthal bins which are illustrated in Fig. 4a. The bins (blue) along the image horizontal are aligned with projections of the strain tensor $\bm{\varepsilon}$ near the $\bm{x}$ direction ($\varepsilon_{xx}$), while vertical bins (red) are projections near the $\bm{z}$ direction ($\varepsilon_{zz}$). The other four sets of bins (orange, green, teal, and purple) further capture the anisotropy of the diffraction ring evolution. Once integrated, the six sets of 1D diffraction profiles were concatenated into a single vector illustrated in 4b. This choice of binning strategy was made as a compromise between probing a sufficient number of grains along different directions while still capturing the anisotropy of diffraction ring shape changes due to multiaxial elastic strains.
Figure 4: a) Illustration of the six sets of azimuthal bins (colored) around which diffraction images were integrated. b) Concatenation of the 1D intensity data from the different bins for use in ML models. A comparison of representative experimental and simulated diffraction data are shown.
@@ 3.1 Heat Transfer and Fluid Flow
Specific Heat ($\mathrm{J}\mathrm{/}\mathrm{k}\mathrm{g}\mathrm{/}\mathrm{K}$) $360.4+0.26~{}T-4\times 10^{-6}~{}T^{2}$
@@
A series of simulations were performed rastering a simulated laser heat source over the top of wall specimens with the same thickness (0.53 mm) and height (3 mm) as in the experiment. The length of the specimen was 11.5 mm to reduce computational cost. The heat source moved along the top surface (normal $\bm{z}$) in the $\bm{x}$ direction. Following [20], nine different simulations were performed around the nominal experimental laser setting (power and speed varied) with the same approximate laser size and are summarized in Table 2. The extra thermomechanical simulations beyond 120 W and 0.04 m/s help to compensate for any uncertainty in the laser parameters applied during the experiment. After the heat source passed over the length of the specimen, the sample was allowed to cool for 5 s as mentioned. The primary outputs are temperature fields $T(\bm{r})$, where $\bm{r}$ denotes position, saved at a rate of 500 Hz as the heat source moved and then 100 Hz as the sample cooled for a total of 150 time steps. As mentioned, the code adaptively changes the calculation grid throughout the simulation for computational efficiency. To prepare the temperature fields for use in the next steps of the workflow (Fig. 1), the simulations were post-processed using bilinear interpolation to remap the output to a regular grid with a point spacing of 20 $\mu$m.
@@ 3.2 Thermomechanical Model
Once the 9 time series temperature fields were calculated from the heat transfer and fluid flow modeling, they are used as external data inputs for an elasto-plasticity model in ANSYS to generate thermal $\varepsilon^{T}(\bm{r})$ and elastic strain $\bm{\varepsilon}^{E}(\bm{r})$ fields. The element type used was eight node ‘brick’ elements with three degrees of freedom (translations in $\bm{x}$, $\bm{y}$, and $\bm{z}$) at each node. Element sizes ranged from 125 to 140 $\mu$m (approximately 10,000 elements per simulations) along their edges.
@@
where $\sigma^{Y}(T)$ is the initial temperature-dependent yield strength, $H(T)$ is the temperature-dependent hardening rate, $\tilde{\varepsilon}^{P}$ is the equivalent plastic strain, and $\tilde{\sigma}$ is the equivalent (von Mises) stress. If the stresses are sufficient to initiate yielding (and ultimately the development of residual stress), plastic flow occurs. As the literature is limited regarding temperature-dependent properties of specific alloy compositions, the temperature-dependent coefficient of thermal expansion $\alpha(T)$ and the elastic moduli, Young’s modulus $E(T)$ and Poisson’s ratio $\nu(T)$, were determined using first-principles DFT. The process by which these values were calculated are presented in Appendix A1. The thermal expansion and elastic moduli used for thermomechanical model input are provided in Fig. 5. Temperature-dependent coefficient of thermal expansion measured from the same IN625 build as the experimental thin wall with dilatometry and used to evaluate the DFT results are also provided in Fig. 5a. The tabulated yield and hardening parameters were determined from a combination of mechanical tests ($\leq 500^{\circ}$) performed on the same material and IN625 parameters built into the ANSYS package (Table 3).
Figure 5: Temperature-dependent a) coefficient of thermal expansion $\alpha$, b) isotropic Young’s modulus $E$, and c) isotropic Poisson’s ratio $\nu$ used as input for the finite element simulations of the development of thermal and residual stresses in the AM IN625 wall specimens derived from the single crystal elastic moduli calculated from DFT. Dilatometer measurements of thermal expansion used to evaluate the DFT results are also provided in a).
Each time step takes the previously calculated thermal field and calculates local thermal strains (Eq. 2) from the temperature-dependent coefficient of thermal expansion previously calculated using DFT [28]. A strain field is then calculated which satisfies both compatibility (in conjunction with the thermal and plastic strain) and mechanical equilibrium. In all thermomechanical simulations, the thermal gradient was sufficient to initiate yielding and plastic flow. While plastic strain does not directly alter the crystal lattice and resulting diffraction, the deformation incompatibility created by the plastic strain gives rise to elastic strain and stress distributions which are measurable through diffraction. Fig. 6 shows representative fields ($P=120$ W, $v=0.05$ m/s, and spot diameter of 100 $\mu$m) calculated by the thermomechanical model as the laser passes over the sample (Figs. 6a, 6c, 6e, and 6g) and after cooling (Figs. 6b, 6d, 6f, and 6h). The thermal and elastic strains are plotted with different color scales due to the large differences in maximum magnitude. Important to note are the relatively large residual tensile $\varepsilon^{E}_{xx}$ strains that develop during the cooling process.
@@
Figure 6: Representative fields from the thermomechanical simulations ($P=120$ W, $v=0.05$ m/s, and spot diameter of 100 $\mu$m) as the laser is passing overhead: a) $T$, c) $\varepsilon^{T}$, e) $\varepsilon^{E}_{xx}$, and g) $\varepsilon^{E}_{zz}$. Representative fields from the same laser parameter simulation after cooling: b) $T$, d) $\varepsilon^{T}$, f) $\varepsilon^{E}_{xx}$, and h) $\varepsilon^{E}_{zz}$.
To facilitate X-ray diffraction simulations from the elasto-plasticity results, a series of scripts were developed to convert ANSYS mechanical data into a format compatible with X-ray simulations. This effort included developing (i) a new Python based binding for ASCII output of ANSYS data; (ii) a script to convert element position, stress, and strain data into the coordinate system used by HEXRD; and (iii) developing a framework to interpolate thermomechanical data on to a finer grid (20 $\mu$m spacing) necessary for diffraction simulations.